Page Turner loves discrete mathematics. She has 7 "graph theory" books, 8 books about combinatorics, and 7 "set theory" books. How many ways can she place her discrete mathematics books on the same shelf in a row if:
a) there are no restrictions.
b) graph theory books are next to each other but the others could be anywhere on the shelf.

c) books are organized by their topic (same kinds are next to each other).

Answer by sudhanshu_kmr(1152)


7 graph theory, 8 combinatorics and 7 set theory books

a) 7+8+7 = 22
ans: 22!
b) 1 + 8 + 7 = 16
ans : 7! * 16!

c) 1 + 1 + 1 = 3
ans: 7! * 8! * 7! * 3!

If any doubt, feel free to contact ...